OpenStudy (anonymous):

In an election, only R, S and T received votes. If S received twice as many votes as R, and T received three times as many votes as S, what percent of the total vote was received by T?

OpenStudy (anonymous):

vote R has = x vote S has = 2 R = 2x Vote T has = 3 S = 3 *(2x) = 6x total vote = 1x + 2x + 6x = 9x = 100% of the votes find x value by100% / 9 = 11.11 % votes Vote received by T = 6x = 6 * (11.11%) = 66.66%

OpenStudy (anonymous):

Thank you!